OpenStudy (kc_kennylau):

Find the smallest positive integer x such that x≡0 (mod 16) and x≡2 (mod 25).

OpenStudy (kc_kennylau):

They are in 1 question. I'll rephrase it.

OpenStudy (anonymous):

for this i would grind it till you find it

OpenStudy (inkyvoyd):

ooh, booleon AND sorry

OpenStudy (kc_kennylau):

I want to use Chinese Remainder Theorem, but I don't know how to use it.

OpenStudy (anonymous):

16, 32, 48, ... one has to be 2 mod 25

OpenStudy (kc_kennylau):

I want to use Chinese Remainder Theorem, but I don't know how to use it.

OpenStudy (inkyvoyd):

I can look at my book and try to work it out, but it'll take a little while. We never covered chiense remainder theorem in discrete class since we ran outta time

OpenStudy (kc_kennylau):

@satellite73 ?

OpenStudy (anonymous):

there is an actual algorithm for finding this you can start with \(x=16t\) then \[16t\equiv 2 (25)\] but at this step it is probably easiest to keep adding 16 until you get it

OpenStudy (kc_kennylau):

I want to explicitly use the Chinese Remainder Theorem.

OpenStudy (inkyvoyd):

ok, my book says to construct solution of x≡a_1(mod m_1) x≡a_2(mod m_2) x≡a_3(mod m_3) by letting x=a_1M_1y_1+a_2M_2y_2+...+a_n*M_n*y_n where M_n equals the product of all moduli except m_n and M_ky_k≡1(mod m_k)

OpenStudy (kc_kennylau):

what is y?

OpenStudy (anonymous):

take a look at this example your is going to be easier because you are only solving 2 http://www.youtube.com/watch?v=3PkxN_r9up8

OpenStudy (inkyvoyd):

I'm not providing the reasoning of what the solution is simultaenous, but I'll just try to work out the problem x≡0 (mod 16) x≡2(mod 25) since the two are pairwise relatively prime, we can continue M_1=25 M_2=16 a_1=0 a_2=2 we know 25y_1≡1(mod 16) and we know that 16y_2≡1(mod 25)

OpenStudy (inkyvoyd):

in other words y_1=9 and y_2=11

OpenStudy (inkyvoyd):

thus the solution is x=a_1M_1y_1+a_2M_2y_2, or x=0+2*16*11=352

OpenStudy (inkyvoyd):

@kc_kennylau , does this make sense?

OpenStudy (kc_kennylau):

Thanks :)

OpenStudy (kc_kennylau):

but how do you know that y_2=11?

OpenStudy (inkyvoyd):

M_ky_k≡1(mod m_k)

OpenStudy (kc_kennylau):

brute force? :/

OpenStudy (inkyvoyd):

\(\large M_kY_k≡1(mod m_k)\)

OpenStudy (inkyvoyd):

so yes, I solved 16y_2≡1(mod 25) by brute force, though I am sure there is an easier way to do it.

OpenStudy (kc_kennylau):

but there is no easier way to do it :P

OpenStudy (kc_kennylau):

anyway thanks :D

OpenStudy (inkyvoyd):

np - looked around and I think the solution to a linear congruence is beyond the scope of my DM course (and perhaps what you are learning)